OpenStudy (anonymous):

can someone explain to me how to solve for x in \frac{1}{x-1}-\frac{1}{2x+1}=0

hartnn (hartnn):

add 1/(2x+1) to both sides.

hartnn (hartnn):

\(\huge \frac{1}{x-1}-\frac{1}{2x+1}+\frac{1}{2x+1}=0+\frac{1}{2x+1} \) what remains on left side ?

OpenStudy (anonymous):

\[\frac{ 1 }{ x-1? } remains on the \left\]

jhonyy9 (jhonyy9):

nursingmom2 you need to cooperate anything not just gett the right answers ok ?

hartnn (hartnn):

yes, thats correct. so u have \(\huge \frac{1}{x-1}=\frac{1}{2x+1}\) now i will give u hint : if \(\huge \frac{1}{a}=\frac{1}{b} \quad \text {then a=b}\) can u simplify further ?

OpenStudy (anonymous):

yes, if you have x-1=2x+1, you could subtract 1 from both sides giving you x-2=2x, then further simplify by subtracting x from both sides which would give you -2=x

hartnn (hartnn):

you are absolutely correct! good work :)

OpenStudy (anonymous):

thank you

hartnn (hartnn):

welcome :)